Solve.

4x squared = 25

Answer by checkley71(8403)

4X^2=25
4X^2-25=0
(2X+5)(2X-5)=0
2X+5=0
2X=-5
X=-5/2 OR -2.5 ANSWER.
2X-5=0
2X=5
X=5/2 OR 2.5 ANSWER.